# GUARD: Guided Unlearning and Retention via Data Attribution for Large Language Models

### v1

@@ 6 Conclusion
Limitations and future work. While GUARD significantly improves existing baselines, there remains room for potential advances. Future work may explore a) integrating noisy unlearning techniques [5] for added robustness; b) adopting more sophisticated data attribution methods; and c) developing mechanisms to handle sequential unlearning requests. These directions can potentially enhance GUARD’s adaptability and effectiveness in more complex, large-scale unlearning scenarios.
Broader impact. LLM unlearning promotes privacy by enabling selective data removal. GUARD highlights the potential of attribution-guided methods for safer, more reliable unlearning, highlighting challenges in balancing unlearning quality with model utility.
@@
Appendix A Related Works
A.1 Unlearning for LLM
Existing LLM Unlearning approaches broadly fall into two categories: fine-tuning-based and inference-time-based methods. Finetuning approaches such as Gradient Ascent [50], Preference Optimization [29] and variants thereof, modify model weights by adjusting training objectives, typically through loss maximization on the forget set [53, 20] and incorporation of regularization terms based on KL divergence, retained data consistency, or prompt-guided responses [40, 6, 46]. Recent approaches such as WAGLE [15] improve computational efficiency by isolating the parameters most relevant to forgetting. However, fine-tuning-based methods frequently suffer from excessive forgetting and retention degradation [54], compromising the utility of the retained model knowledge. A complementary line of work explores inference-time unlearning [35, 12, 41], which avoids model updates by modifying prompts or logits at inference to suppress specific outputs [27, 14]. While computationally attractive, these methods do not guarantee actual forgetting, as sensitive information remains embedded in the unchanged model parameters. Emerging directions [52, 55] also investigate properties beyond forgetting and retention, such as robustness of unlearning mechanisms and recovery from catastrophic forgetting.
Fine-tuning-Based Methods. Many unlearning techniques modify model weights through loss-based fine-tuning strategies, such as gradient ascent (GA) [50], preference optimization (PO) [29], negative preference optimization (NPO) [53], and representation misdirection [20]. These methods typically maximize the loss on the forget set while incorporating regularization objectives such as KL divergence minimization, joint training with retain data [40], or prompt-based desirable loss surrogates [6]. Despite their effectiveness in forgetting, these methods often degrade performance on retained knowledge due to excessive forgetting [54]. WAGLE [15] focuses on improving computation efficiency by isolating critical parameters for forgetting via weight attribution, whereas [46] adds a regularization term to penalize low-quality generations, yet both empirically fail to preserve retention. Other fine-tuning efforts include [8, 34, 16], but none achieve robust retention preservation.
@@ A.1 Unlearning for LLM
Input-level methods operate at inference time by manipulating inputs to suppress undesired outputs without altering model weights. These include prompt filtering [27], instruction-based tuning [35], and logit-based interventions [12, 14]. ECO Prompts [27] use prompt classifiers and zeroth-order optimization to induce forgetting, while Unlearning from Logit Difference (ULD) [14] employs an assistant LLM to reduce side effects. Although computationally efficient, these methods rely on access to sensitive data at inference time, limiting their privacy guarantees [41, 26].
@@
Recent efforts explore complementary directions beyond weight-based or input-level interventions. Multimodal unlearning [19], adversarial robustness [38], and standardized benchmarks [52, 44] enhance evaluation protocols but do not directly address retention. Utility-guided methods, such as gradient-based influence modeling [45], propose retention-aware metrics without offering actionable strategies. Other approaches like quantization-based recovery [55] attempt to restore forgotten knowledge post hoc but remain orthogonal to the goal of preserving retention during unlearning.
A.2 Data Attribution
Data attribution in machine learning aims to quantify the influence of individual data points on model behavior, supporting tasks such as debugging, data cleaning, and interpretability. Classical approaches, such as influence functions, estimate the effect of removing a single data point on model parameters or predictions. Koh and Liang [17] adapted influence functions to deep learning, illustrating their value in identifying mislabeled data and interpreting model decisions. However, subsequent studies revealed notable limitations. Basu et al. [2] showed that influence estimates are highly sensitive to hyperparameters and model architecture, particularly in deep networks. Han and Liu [10] further found that these methods can fail basic sanity checks, attributing high influence to unrelated data. Bae et al. [1] also demonstrated that influence functions may estimate unintended quantities in deep neural networks, raising concerns about their reliability in complex models.
To address the limitations of classical influence functions, recent work has introduced more robust and scalable data attribution techniques. Among these, direct estimators such as regression-based datamodels have demonstrated improved predictive accuracy and generalization, particularly in large-scale settings where computational resources are sufficient [9]. Two state-of-the-art approaches—TRAK [9] and EK-FAC [13]—build upon this paradigm and offer practical solutions across diverse modalities. TRAK approximates a deep neural network with a linear surrogate model and estimates influence by ensembling over multiple model checkpoints and random projections. EK-FAC improves efficiency through a combination of strategies: partitioning the Hessian into block-diagonal components, expressing gradients as Kronecker products of smaller factors, and applying spectral regularization to stabilize curvature estimates.
@@ A.2 Data Attribution
Appendix B Theoretical soundness of our data attribution tailored for LLM Unlearning
B.1 Preliminiaries: conventional data attribution
@@
Data attribution quantifies the impact of a specific training datapoint on the model’s loss under the Leave-One-Out framework [36]. Given a dataset $D$, the influence of a datapoint $(x_{j},y_{j})$ on the model’s optimal loss is defined as:
@@
The Influence Function [1] provides an efficient approximation for the leave-one-out influence. Let $H_{D}(\theta)$ denote the Hessian matrix summing over data-wise second derivatives: $H_{D}(\theta):=\sum_{(x_{i},y_{i})\in D}\nabla^{2}l(x_{i},y_{i};\theta),$ The influence function for a datapoint $(x_{j},y_{j})$ is then given by:
@@ B.2 GUARD attribution as bounds
Based on these notations, we propose Assumptions B.1 , under which we derive a bound for $a_{j}^{\text{IF}}$ as defined in Equation 7, as stated in Lemma B.1
Assumption B.1 (Smoothness of the Loss Function).
@@
Remark.
Many commonly used loss functions, such as squared loss and logistic loss, satisfy this assumption. This assumption ensures that the Hessian matrix exists and is well-behaved, allowing for stable second-order optimization techniques such as Newton’s method.
Lemma B.1 (Bounds on the Influence Function by Proxy Attribution).
Suppose that Assumptions B.1 holds. Define $q^{+}:=\sum_{k}[\bm{1}_{q_{k}\geq 0}\cdot q_{k}]$, and $q^{-}:=\sum_{k}[\bm{1}_{q_{k}<0}\cdot q_{k}]$. The influence function estimator $a_{j}^{\text{IF}}$, as defined in Equation equation 7, satisfies the following bound:
@@
Remark.
The proof is provided below. As established in Lemma B.1, the influence function $a_{j}^{\text{IF}}$ is scaled by the GUARD proxy attribution $a_{j}$. Building on this intuition, we later approximate $a_{j}^{\text{IF}}$ by $C\cdot a_{j}^{\text{GUARD}}$, where $C$ is a constant that will be canceled by unification (See Section 3.1). A discussion on $C$ is deferred to Appendix B.3. Note that the scale constant $C$ is cancelled by obtaining GUARD multiplier using Eq. 4. Therefore, in practical implementation, we adopt the unscaled proxy attribution:
@@
As $H_{D}(\theta_{0})$ is symmetric, we have $H_{D}(\theta_{0})^{-1}$ is asymmetric, therefore, the eigenvectors of $H_{D}(\theta_{0})^{-1}$ are orthogonal to each other, i.e., $v_{i}\cdot v_{j}=0,\forall i\neq j$. Therefore,
@@
which is the upper bound in Lemma B.1 Similarly,
@@
which is the lower bound in Lemma B.1
B.3 Discussion on $C$
By Lemma B.1, the influence function estimate $a_{j}^{\text{IF}}$ is scaled by the GUARD proxy attribution $a_{j}^{GUARD}$, satisfying the bound:
@@ B.3 Discussion on $C$
Approximating $\hat{\Delta}_{L}^{j}$ by a globally optimized scaling factor $C$, we analyze the approximation error:
@@
where $\overline{C}_{GUARD}:=\frac{\sum_{j\in I_{f}}C_{j}\cdot[a_{j}^{GUARD}]^{2}}{\sum_{j\in I_{f}}[a_{j}^{GUARD}]^{2}}$ is the average of $\{C_{j}\}$ weighted by $[a_{j}^{GUARD}]^{2}$.
Appendix C Proof for theoretical guarantee
Lemma C.1 (Restatement of Lemma 4.1).
Under Assumption 4.1, the loss on the retain set under GUARD is lower than that under GA by:
@@ Appendix C Proof for theoretical guarantee
Based on loss definitions and model update rules, we get the approximations of the increase in the retain loss compared with the model before unlearning. Especially, the increase in the loss on the retain set induced by GA is given by
@@
$=\eta\cdot\kappa+{\mathcal{O}}(\delta_{\theta}^{2})$ (11)
@@
where $\overline{g}^{\prime}_{f}:=\frac{1}{n_{f}}\sum_{i\in I_{f}}\left[\omega_{i}^{\text{GUARD}}\cdot g_{i}\right]$. Denote $\kappa^{\prime}:=\langle\overline{g}_{r},\overline{g}^{\prime}_{f}\rangle$. It holds that
@@
With slight abuse of notation, define the alignment between averaged gradient of the retain set and datawise gradient of the forget set as $\kappa_{j}:=\langle\overline{g}_{r},g_{j}\rangle,\ j\in I_{f}$. Define its variance as $\sigma_{\kappa}^{2}:=\frac{1}{n_{f}}\sum_{j\in I_{f}}[\kappa_{j}]^{2}-[\frac{1}{n_{f}}\sum_{j\in I_{f}}\kappa_{j}]^{2}$, equivalent to $\sum_{j\in I_{f}}{\kappa_{j}}^{2}=n_{f}\cdot(\sigma^{2}_{\kappa}+\kappa^{2})$. Therefore,
@@
Lemma C.2 (Restatement of Lemma 4.2).
Under Assumption 4.1, the difference in forget loss between GUARD and GA is bounded by:
@@
Suppose the second and higher orders of the norm of change in model weights can be omitted in the theoretical analysis, then based on loss definitions and model update rules, we get the approximations of the increase in the forget loss compared with the model before unlearning. Especially, the increase in the loss on the forget set induced by GA is given by
@@
where $\overline{g}^{\prime}_{f}$ is the weighted average gradient over forget set, $\overline{g}^{\prime}_{f}:=\frac{1}{n_{f}}\sum_{i\in I_{f}}[\omega_{i}^{\text{GUARD}}\cdot g_{i}]$.
@@
$=\frac{\tau}{{n_{f}}(\tau-\kappa)}\left[{n_{f}}\|\overline{g}_{f}\|_{2}^{2}-\frac{1}{\tau}\sum_{j\in I_{f}}\left[\overline{g}_{r}^{\top}\cdot g_{j}\cdot g_{j}^{\top}\cdot\overline{g}_{f}\right]\right]+{\mathcal{O}}(\delta^{2}_{\kappa})$ (16)
Let $d$ be the dimensionality of the gradient vectors. Then we can rewrite:
@@
Denote $\text{Cov}_{f}$ as the sample covariance matrix of $g_{j}$, $j\in I_{f}$, defined as $\text{Cov}_{f}:=\frac{1}{n_{f}}\sum_{j\in I_{f}}(g_{j}-\overline{g}_{f})(g_{j}-\overline{g}_{f})^{\top}$. Then we have
@@
$=(1+\delta_{\kappa})\|\overline{g}_{f}\|_{2}^{2}+{\mathcal{O}}(\delta_{\kappa}^{2})$ (19)
@@
Under Assumption 4.1, GUARD reduces the Sacrifice Rate relative to GA as:
@@
C.1 Key factors analysis for theoretical guarantee

### v2

@@ 6 Conclusion
Ethics Statement
The authors acknowledge adherence to the ICLR Code of Ethics and recognize that this work addresses legitimate privacy concerns by developing methods to selectively remove sensitive information from large language models in support of data protection regulations. We honestly report that current unlearning methods, including GUARD, may not provide absolute guarantees of information removal, and practitioners should carefully evaluate models for unintended bias or fairness implications after unlearning procedures. All experiments were conducted using established benchmarks with proper attribution and transparent reporting of results and limitations. This research contributes to making AI systems more controllable and privacy-preserving, though we recognize that responsible deployment requires interdisciplinary collaboration with legal, policy, and ethics experts. The authors commit to sharing code and experimental details to support reproducible research while being mindful of potential misuse scenarios.
Reproducibility Statement
To ensure reproducibility of our results, we have made comprehensive efforts to document all experimental details and follow established evaluation protocols. Our experiments strictly adhere to the evaluation frameworks and protocols specified in the original TOFU (Maini et al., 2024b) and MUSE (Shi et al., 2024a) benchmark papers, and we encourage researchers to follow the guidance provided in those publications for proper reproduction of baseline comparisons. All hyperparameter settings, model configurations, and experimental setup details for the GUARD method are provided in Appendix F.1, including learning rates, batch sizes, temperature parameters, and hardware specifications used across different model architectures. The algorithmic implementation of GUARD is detailed in Algorithm 1 and Algorithm 2, with mathematical formulations provided in Section 3 and theoretical derivations included in Appendix D. Our evaluation metrics and experimental protocols for both absolute performance and sacrifice rate measurements are comprehensively described in Section 5, with additional results across different data splits and model architectures provided in Appendices F.2 through F.4. We commit to releasing our implementation code and detailed experimental scripts to facilitate exact reproduction of our reported results.
@@
Appendix A LLM Usage Statement
We used a large language model (LLM) Claude Sonet 4 for language polishing of parts of the manuscript (phrasing, grammar, and clarity). The LLM was only used to improve presentation; it did not contribute to the technical content, experimental design, data analysis, or conclusions. All substantive scientific claims, results, and interpretations are the authors’ own, and the authors take full responsibility for the final text. Any LLM-generated text was reviewed, edited, and approved by the authors.
Appendix B Related Works
B.1 Unlearning for LLM
Existing LLM Unlearning approaches broadly fall into two categories: Fine-tuning-based and inference-time-based methods. Fine-tuning approaches such as Gradient Ascent (GA) Yao & et al. (2023), Preference Optimization Maini & et al. (2024) and variants thereof, modify model weights by adjusting training objectives, typically through loss maximization on the forget set Zhang & et al. (2024a); Li & et al. (2024) and incorporation of regularization terms based on KL divergence, retained data consistency, or prompt-guided responses Shi & et al. (2024); Choi & et al. (2024); Wang et al. (2024a). Recent approaches such as WAGLE Jia et al. (2024) improve computational efficiency by isolating the parameters most relevant to forgetting. However, fine-tuning-based methods frequently suffer from excessive forgetting and retention degradation Zhang & et al. (2024b), compromising the utility of the retained model knowledge. A complementary line of work explores inference-time unlearning Pawelczyk & et al. (2023); Huang & et al. (2024); Thaker & et al. (2024), which avoids model updates by modifying prompts or logits at inference to suppress specific outputs Liu et al. (2024); Ji et al. (2024). While computationally attractive, these methods do not guarantee actual forgetting, as sensitive information remains embedded in the unchanged model parameters. Emerging directions Yao et al. (2024b); Zhang et al. (2024b) also investigate properties beyond forgetting and retention, such as robustness of unlearning mechanisms and recovery from catastrophic forgetting.
Fine-Tuning-Based Methods. Many unlearning techniques modify model weights through loss-based fine-tuning strategies, such as GA Yao & et al. (2023), preference optimization (PO) Maini & et al. (2024), negative preference optimization (NPO) Zhang & et al. (2024a), and representation misdirection Li & et al. (2024). These methods typically maximize the loss on the forget set while incorporating regularization objectives such as KL divergence minimization, joint training with retain data Shi & et al. (2024), or prompt-based desirable loss surrogates Choi & et al. (2024). Despite their effectiveness in forgetting, these methods often degrade performance on retained knowledge due to excessive forgetting Zhang & et al. (2024b). WAGLE Jia et al. (2024) focuses on improving computation efficiency by isolating critical parameters for forgetting via weight attribution, whereas Wang et al. (2024a) adds a regularization term to penalize low-quality generations, yet both empirically fail to preserve retention. Other fine-tuning efforts include Eldan & Russinovich (2023); Patil & et al. (2024); Jia & et al. (2024), but none achieve robust retention preservation.
@@ B.1 Unlearning for LLM
Input-level methods operate at inference time by manipulating inputs to suppress undesired outputs without altering model weights. These include prompt filtering Liu et al. (2024), instruction-based tuning Pawelczyk & et al. (2023), and logit-based interventions Huang & et al. (2024); Ji et al. (2024). ECO Prompts Liu et al. (2024) use prompt classifiers and zeroth-order optimization to induce forgetting, while Unlearning from Logit Difference (ULD) Ji et al. (2024) employs an assistant LLM to reduce side effects. Although computationally efficient, these methods rely on access to sensitive data at inference time, which limits their privacy guarantees Thaker & et al. (2024); Liu & et al. (2024).
@@
Recent efforts explore complementary directions beyond weight-based or input-level interventions. Multimodal unlearning Li et al. (2024a), adversarial robustness Schwinn et al. (2024), and standardized benchmarks Yao et al. (2024b); Wang et al. (2025a) enhance evaluation protocols but do not directly address retention. FLAT Wang et al. (2024b) eliminates the need for Retain Data by maximizing the f-divergence between template and forget responses, while Yang et al. Yang et al. (2025) analyze loss reweighting mechanisms through saturation and importance-based categorizations. Negative Preference Optimization (NPO) Zhang et al. (2024a) treats Forget Data as negative responses in preference optimization, providing bounded objectives with adaptive gradient smoothing to prevent catastrophic collapse during unlearning. SimNPO Fan et al. (2025) further addresses reference model bias in NPO by adopting a reference-free, length-normalized approach that more effectively allocates optimization power across forget samples with varying difficulty levels. Utility-guided methods, such as gradient-based influence modeling Wang et al. (2025b), propose retention-aware metrics without offering actionable strategies. Other approaches like quantization-based recovery Zhang et al. (2024b) attempt to restore forgotten knowledge post hoc but remain orthogonal to the goal of preserving retention during unlearning.
B.2 Data Attribution
Data attribution in machine learning aims to quantify the influence of individual data points on model behavior, supporting tasks such as debugging, data cleaning, and interpretability. Classical approaches, such as influence functions, estimate the effect of removing a single data point on model parameters or predictions. Koh & Liang (2017) adapted influence functions to deep learning, illustrating their value in identifying mislabeled data and interpreting model decisions. However, subsequent studies revealed notable limitations. For example,  Basu et al. (2021) showed that influence estimates are highly sensitive to hyperparameters and model architecture, particularly in deep networks. Han & Liu (2022) further found that these methods can fail basic sanity checks, attributing high influence to unrelated data. Bae et al. (2022) also demonstrated that influence functions may estimate unintended quantities in deep neural networks, raising concerns about their reliability in complex models.
To address the limitations of classical influence functions, recent work has introduced more robust and scalable data attribution techniques. Among these, direct estimators such as regression-based datamodels have demonstrated improved predictive accuracy and generalization, particularly in large-scale settings where computational resources are sufficient Ellis et al. (2022). Two state-of-the-art approaches—TRAK Ellis et al. (2022) and EK-FAC Ingram et al. (2023)—build upon this paradigm and offer practical solutions across diverse modalities. TRAK approximates a deep neural network with a linear surrogate model and estimates influence by ensembling over multiple model checkpoints and random projections. EK-FAC improves efficiency through a combination of strategies: partitioning the Hessian into block-diagonal components, expressing gradients as Kronecker products of smaller factors, and applying spectral regularization to stabilize curvature estimates.
@@ B.2 Data Attribution
Appendix C Theoretical soundness of our data attribution tailored for LLM Unlearning
C.1 Preliminiaries: conventional data attribution
@@
Data attribution quantifies the impact of a specific training datapoint on the model’s loss under the Leave-One-Out framework Rad & Maleki (2018). Given a dataset $D$, the influence of a datapoint $(x_{j},y_{j})$ on the model’s optimal loss is defined as:
@@
The Influence Function Bae et al. (2022) provides an efficient approximation for the leave-one-out influence. Let $H_{D}(\theta)$ denote the Hessian matrix summing over data-wise second derivatives: $H_{D}(\theta):=\sum_{(x_{i},y_{i})\in D}\nabla^{2}l(x_{i},y_{i};\theta),$ The influence function for a datapoint $(x_{j},y_{j})$ is then given by:
@@ C.2 GUARD attribution
Based on these notations, we propose Assumptions 4 , under which we derive a bound for $a_{j}^{\text{IF}}$ as defined in Eq. 6, as stated in Lemma 4
Assumption 4 (Smoothness of the Loss Function).
@@
Remark 1.
Many commonly used loss functions, such as the squared loss and logistic loss, satisfy this assumption. This assumption ensures that the Hessian matrix exists and is well-behaved, allowing for stable second-order optimization techniques such as Newton’s method.
Lemma 4 (Bounds on the Influence Function by Proxy Attribution).
Suppose that Assumptions 4 holds. Define $q^{+}:=\sum_{k}[\bm{1}_{q_{k}\geq 0}\cdot q_{k}]$, and $q^{-}:=\sum_{k}[\bm{1}_{q_{k}<0}\cdot q_{k}]$. The influence function estimator $a_{j}^{\text{IF}}$, as defined in Eq. 6, satisfies the following bound:
@@
Remark 2.
The proof is provided below. As established in Lemma 4, the influence function $a_{j}^{\text{IF}}$ is scaled by the GUARD proxy attribution $a_{j}$. Building on this intuition, we later approximate $a_{j}^{\text{IF}}$ by $C\cdot a_{j}^{\text{GUARD}}$, where $C$ is a constant that will be canceled by unification (See Section 3.1). A discussion on $C$ is deferred to Appendix C.3. Note that the scale constant $C$ is cancelled by obtaining GUARD multiplier using Eq. 3. Therefore, in practical implementation, we adopt the unscaled proxy attribution:
@@
As $H_{D}(\theta_{0})$ is symmetric, we have that $H_{D}(\theta_{0})^{-1}$ is asymmetric, and therefore, the eigenvectors of $H_{D}(\theta_{0})^{-1}$ are orthogonal to each other, i.e., $v_{i}\cdot v_{j}=0,\forall i\neq j$. Consequently,
@@
which is the upper bound in Lemma 4. Similarly,
@@
which is the lower bound in Lemma 4
C.3 Choosing the scaling factor $C$
By Lemma 4, the influence function estimate $a_{j}^{\text{IF}}$ is scaled by the GUARD proxy attribution $a_{j}^{GUARD}$, satisfying the bound:
@@ C.3 Choosing the scaling factor $C$
Approximating $\hat{\Delta}_{L}^{j}$ by a globally optimized scaling factor $C$, we obtain the following approximation error
@@
where $\overline{C}_{GUARD}:=\frac{\sum_{j\in I_{f}}C_{j}\cdot[a_{j}^{GUARD}]^{2}}{\sum_{j\in I_{f}}[a_{j}^{GUARD}]^{2}}$ is the average of $\{C_{j}\}$ weighted by $a_{j}^{GUARD}]^{2}$.
Appendix D Proofs of theoretical guarantees
Lemma 5 (Restatement of Lemma 1).
Under Assumption 1, the loss on the retain set under GUARD is lower than that under GA by:
@@ Appendix D Proofs of theoretical guarantees
Based on the loss definitions and model update rules, we get the approximations of the increase in the retain loss compared with the model before unlearning. Specifically, the increase in the loss on the retain set induced by GA is given by
@@
$=\eta\cdot\kappa+{\mathcal{O}}(\delta_{\theta}^{2}),$ (10)
@@
where $\overline{g}^{\prime}_{f}:=\frac{1}{n_{f}}\sum_{i\in I_{f}}\left[\omega_{i}^{\text{GUARD}}\cdot g_{i}\right]$. Let $\kappa^{\prime}:=\langle\overline{g}_{r},\overline{g}^{\prime}_{f}\rangle$. It holds that
@@
With a slight abuse of notation, define the alignment between averaged gradients of the Retain Set and datawise gradients of the Forget Set as $\kappa_{j}:=\langle\overline{g}_{r},g_{j}\rangle,\ j\in I_{f}$. Define its variance as $\sigma_{\kappa}^{2}:=\frac{1}{n_{f}}\sum_{j\in I_{f}}[\kappa_{j}]^{2}-[\frac{1}{n_{f}}\sum_{j\in I_{f}}\kappa_{j}]^{2}$, equivalent to $\sum_{j\in I_{f}}{\kappa_{j}}^{2}=n_{f}\cdot(\sigma^{2}_{\kappa}+\kappa^{2})$. Therefore,
@@
Lemma 6 (Restatement of Lemma 2).
Under Assumption 1, the difference in the forget loss between GUARD and GA is bounded as
@@
Suppose that the second and higher orders of the norm of change in model weights can be omitted in the theoretical analysis. Then, based on loss definitions and model update rules, we get the approximations of the increase in the forget loss compared with the model before unlearning. Specifically, the increase in the loss on the forget set induced by GA is given by
@@
where $\overline{g}^{\prime}_{f}$ is the weighted average gradient over forget set, $\overline{g}^{\prime}_{f}:=\frac{1}{n_{f}}\sum_{i\in I_{f}}[\omega_{i}^{\text{GUARD}}\cdot g_{i}]$. Also,
@@
$=\frac{\tau}{{n_{f}}(\tau-\kappa)}\left[{n_{f}}\|\overline{g}_{f}\|_{2}^{2}-\frac{1}{\tau}\sum_{j\in I_{f}}\left[\overline{g}_{r}^{\top}\cdot g_{j}\cdot g_{j}^{\top}\cdot\overline{g}_{f}\right]\right]+{\mathcal{O}}(\delta^{2}_{\kappa}).$ (15)
Let $d$ be the dimensionality of the gradient vectors. Then we can write
@@
Let $\text{Cov}_{f}$ be the sample covariance matrix of $g_{j}$, $j\in I_{f}$, defined as $\text{Cov}_{f}:=\frac{1}{n_{f}}\sum_{j\in I_{f}}(g_{j}-\overline{g}_{f})(g_{j}-\overline{g}_{f})^{\top}$. Then we have
@@
$=(1+\delta_{\kappa})\|\overline{g}_{f}\|_{2}^{2}+{\mathcal{O}}(\delta_{\kappa}^{2}).$ (18)
@@
Under Assumption 1, GUARD reduces the Sacrifice Rate relative to GA as:
@@
D.1 Analysis of Key Factors for Theoretical Guarantees